# Isomorphism classes of generating sets

### v1

@@ 4. At a measurable cardinal
This problem was first solved by Džamonja and Shelah [22], where the rough idea is that we force at $\alpha$ with the disjoint sum of all measures which exist in $V[G_{\alpha}]$. This leads to some technical complications in the iteration, which we can avoid using an idea suggested originally by Magidor: the idea is to choose $U_{\alpha}$ using some diamond principle, and organize the construction so that whenever $U_{\alpha}\subseteq U$ (which happens frequently) we have $B_{\alpha}\in U$. Running the iteration for a large enough number of steps, we may produce measures which are generated by arbitrarily long $\subseteq^{*}$-decreasing sequences. The diamond idea appears in a paper by Cummings, Džamonja, Magidor, Morgan and Shelah [18] albeit in a more complex setting.
For technical reasons we will be using “filter Mathias forcing” (see Definition 3.2). If $F$ is a $\kappa$-complete filter, the filter Mathias poset is still $\kappa$-directed closed and $\kappa^{+}$-cc.
Let $\mathbb{Q}$ be a forcing poset and let $\lambda$ be a cardinal. It will be convenient to specify exactly what we mean by a canonical $\mathbb{Q}$-name for a subset of $\lambda$ or a family of subsets of $\lambda$. A canonical $\mathbb{Q}$-name for a subset of $\lambda$ is a name $\dot{\tau}$ such that $\dot{\tau}\subseteq\mathbb{Q}\times\{\check{\alpha}:\alpha<\lambda\}$, and $\{q:(q,\check{\alpha})\in\dot{\tau}\}$ is an antichain for all $\alpha$. A canonical name for a family of subsets of $\lambda$ is a name $\dot{\sigma}$ such that $\dot{\sigma}\subseteq\mathbb{Q}\times\{\dot{\tau}:\mbox{$\dot{\tau}$ is a canonical name for a subset of $\lambda$}\}$, and $\{q:(q,\dot{\tau})\in\dot{\sigma}\}$ is an antichain for all $\dot{\tau}$. It is easy to see that any name for a subset of $\lambda$ or a family of subsets of $\lambda$ is equivalent to a canonical name.
Let $V_{0}\models\mbox{``GCH and $\kappa$ is supercompact''}$”, and let $V=V_{0}[L]$ where $\mathbb{L}$ is the standard Laver preparation to make $\kappa$ indestructible under $\kappa$-directed forcing. Note that GCH holds in $V$ for all cardinals $\delta$ with $\delta\geq\kappa$.
We work in $V$ until further notice. Let $n<\omega$, let $(\lambda_{i})_{0\leq i\leq n}$ be a strictly increasing sequence of regular cardinals with $\kappa<\lambda_{0}$, and let $I=\prod_{i\leq n}\lambda_{i}$. We will define a non-linear iteration $\mathbb{P}$ of ultrafilter Mathias forcing indexed by $I$, arranging that $\mathbb{P}$ is $\kappa$-directed closed and $\kappa^{+}$-cc.
We will also assume that the diamond principle $\diamondsuit_{\lambda_{n}}(\lambda_{n}\cap\operatorname{cof}(\kappa^{+}))$ holds in $V$. This is easy to arrange: prepare $V_{0}$ so that this form of diamond holds there, and then use standard arguments about the preservation of diamond principles by forcing to show that it still holds in $V$ [^1] Our $\mathrm{GCH}$ hypothesis actually implies right away this diamond principle at many, but not all, instances. For example, by Shelah [43] it holds for every successor cardinal $\lambda_{n}=\chi^{+}$ as long as ${\rm cf}\chi\neq\kappa^{+}$. By an unpublished result of the third author, it holds for every $\lambda_{n}$ inaccessible, as $\lambda_{n}$ is above a supercompact cardinal. We will use a form of diamond which is adapted to guessing subsets of $H_{\lambda_{n}}$, which is possible because $|H_{\lambda_{n}}|=\lambda_{n}$. To be precise we will assume that there are sequences $(A_{\alpha})_{\alpha\in\lambda_{n}\cap\operatorname{cof}(\kappa^{+})}$ and $(M_{\alpha})_{\alpha<\lambda_{n}}$ such that
  1. (1)
    $M_{\alpha}\prec H_{\lambda_{n}}$ with $|M_{\alpha}|<\lambda_{n}$.
  2. (2)
    $(M_{\alpha})_{\alpha<\lambda_{n}}$ is continuous and $\bigcup_{\alpha<\lambda_{n}}M_{\alpha}=H_{\lambda_{n}}$.
  3. (3)
    $A_{\alpha}\subseteq M_{\alpha}$, and for every $A\subseteq H_{\lambda_{n}}$ there are stationarily many $\alpha\in\lambda_{n}\cap\operatorname{cof}(\kappa^{+})$ such that $A\cap M_{\alpha}=A_{\alpha}$.
The choice of the sequence $(M_{\alpha})$ is not important, since such sequences always exist and any two such sequences agree on a club subset of $\lambda_{n}$.
As before, we will order $I$ with the product coordinatewise ordering, that is to say $(\eta_{i})_{i\leq n}\leq(\zeta_{i})_{i\leq n}$ if and only if $\eta_{i}\leq\zeta_{i}$ for all $i\leq n$. For $a=(\alpha_{1},\dots,\alpha_{n})\in I$, recall that we denoted
$I_{<a}=\{b\in I:b<a\}=(\alpha_{0}+1)\times...\times(\alpha_{n}+1)\setminus\{a\}.$
As usual ${\mathbb{P}}_{a}$ is defined by recursion on $a$. Part of the definition will involve choosing a ${\mathbb{P}}_{a}$-name $\dot{F}_{a}$ for a $\kappa$-complete uniform filter on $\kappa$ in $V[P_{a}]$, but we postpone the exact definition of $\dot{F}_{a}$ for the moment.
A condition $p\in{\mathbb{P}}_{b}$ is a partial function $p$ such that:
  • $\operatorname{dom}(p)\subseteq I_{<b}$.
  • $|\operatorname{dom}(p)|<\kappa$.
  • For every $a\in\operatorname{dom}(p)$, $p(a)$ is a ${\mathbb{P}}_{a}$-name for a condition in ${\mathbb{Q}}_{a}$, where ${\mathbb{Q}}_{a}$ is the filter Mathias forcing defined in $V^{\mathbb{P}_{a}}$ from the filter $\dot{F}_{a}$.
For conditions $p,q\in{\mathbb{P}}_{b}$, $q\leq p$ if and only if:
  • $\operatorname{dom}(p)\subseteq\operatorname{dom}(q)$.
  • For all $a\in\operatorname{dom}(p)$, $q\restriction I_{<a}\Vdash_{{\mathbb{P}}_{a}}q(a)\leq_{{\mathbb{Q}}_{a}}p(a)$.
The definition makes sense because (by an easy induction) for all $a<b$ we have that ${\mathbb{P}}_{a}$ is a complete subposet of ${\mathbb{P}}_{b}$, and $q\mapsto q\restriction I_{<a}$ is a projection from ${\mathbb{P}}_{b}$ to ${\mathbb{P}}_{a}$. It is easy to verify that ${\mathbb{P}}_{a}$ is $\kappa$-directed closed. For $\kappa^{+}$-cc we first verify that the set of $p$ such that $p\restriction I_{<a}$ determines the stem of $p(a)$ for all $a\in\operatorname{dom}(p)$ is dense, and then use the $\Delta$-system lemma, keeping in mind that $\kappa^{<\kappa}=\kappa$.
Let $\mathbb{P}=\bigcup_{b\in I}{\mathbb{P}}_{b}$, be the direct limit of $\langle\mathbb{P}_{b}\mid b\in I\rangle$. For any downwards closed subset $J$ of $I$, let ${\mathbb{P}_{J}}=\{p\in{\mathbb{P}}:\operatorname{dom}(p)\subseteq J\}$. Then ${\mathbb{P}_{J}}$ is a complete subposet of ${\mathbb{P}}$, and $q\mapsto q\restriction J$ is a projection from ${\mathbb{P}}$ to ${\mathbb{P}_{J}}$ (the proof is identical to Lemma 3.1). In these notations, $\mathbb{P}_{a}=\mathbb{P}_{I_{<a}}$.
We still owe the definition of a ${\mathbb{P}}_{a}$-name $\dot{F}_{a}$ for a $\kappa$-complete uniform filter on $\kappa$. For $\beta<\lambda_{n}$, we say that $\beta$ is an active stage if $\beta\in\lambda_{n}\cap\operatorname{cof}(\kappa^{+})$ and $A_{\beta}$ is a canonical ${\mathbb{P}}_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta}$-name for a measure on $\kappa$. Let $a=(\alpha_{0},\ldots,\alpha_{n-1},\beta)$ and assume for the moment that $\beta$ is an active stage.
Working in $V$, let $F^{0}_{a}$ be the set of canonical ${\mathbb{P}}_{(\alpha_{0}+1)\times...\times(\alpha_{n-1}+1)\times\beta}$-names $\dot{A}$ for subsets of $\kappa$ such that $\Vdash^{V}_{{\mathbb{P}}_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta}}\dot{A}\in A_{\beta}$. This is reasonable because $(\alpha_{0}+1)\times\ldots\times(\alpha_{n-1}+1)\times\beta$ is a downwards closed subset of $\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta$, so that $\dot{A}$ makes sense as a ${\mathbb{P}}_{\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta}$-name for a subset on $\kappa$.
Claim 4.1.
[…]
Proof.
Force to prolong $G_{a}$ to $G$ such that $G\restriction I_{<a}$ is ${\mathbb{P}}_{\downarrow a}$-generic, and let $W=A_{\beta}[G\restriction\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta]$. Then $F^{1}_{a}\subseteq W$ and $V[G\restriction\lambda_{0}\times\ldots\times\lambda_{n-1}\times\beta]\models\mbox{``$W$ is a measure on $\kappa$''}$ ”, and the conclusion follows easily. ∎
If $a=(\alpha_{0},\ldots,\alpha_{n-1},\beta)$ and $\beta$ is an active stage, we let $\mathcal{X}_{a}=\{B_{b}\mid b=(\beta_{1},...,\beta_{n-1},\beta)<a\}$, where $B_{b}$ is the Mathias generic set added by the component $\mathbb{Q}_{b}$ for the filter $F_{b}$. Note that the right-most coordinate is fixed in the definition of $\mathcal{X}_{a}$. If $F^{1}_{a}\cup\mathcal{X}_{a}$ generates a $\kappa$-complete filter, let $F_{a}$ be the $\kappa$-complete uniform filter generated in $V[G_{a}]$ by $F^{1}_{a}\cup\mathcal{X}_{a}$. If $\beta$ is not an active stage or if $F^{1}_{a}\cup\mathcal{X}_{a}$ does not generate a $\kappa$-complete filter, then let $F_{a}$ be the tail filter on $\kappa$.
We record a few easy observations, where we make repeated use of the $\kappa^{+}$-cc and the regularity assumption on the $\lambda_{i}$’s to analyze how subsets of $\kappa$ will appear in the generic extension.
  1. (1)
    Every canonical $\mathbb{P}$-name for a subset of $\kappa$ is a canonical ${\mathbb{P}}_{a}$-name for some $a\in I$. It follows that if $G$ is $\mathbb{P}$-generic then $P(\kappa)\cap V[G]=\bigcup_{a\in I}P(\kappa)\cap V[G_{a}]$, where $G_{a}$ is the projection of $G$ to ${\mathbb{P}}_{a}$.
  2. (2)
    Similarly if $\beta<\lambda_{n}$ and ${\rm cf}(\beta)>\kappa$, then $P(\kappa)\cap V[G_{\prod_{i<n}\lambda_{i}\times\beta}]=\bigcup_{\alpha<\beta}P(\kappa)\cap V[G_{\prod_{i<n}\lambda_{i}\times\alpha}]$.
@@
If $G$ is $\mathbb{P}$-generic then there exist in $V[G]$ a measure $U$ on $\kappa$ and a family of sets $(B_{a})_{a\in I}$ such that
@@
    $B_{a}\in U$ for all $a\in I$.
@@
    For every $Y\in U$ there is $a\in I$ such that $B_{a}\subseteq^{*}Y$.
In particular, $U$ is a simple $P_{I}$-point.
Proof of Theorem 4.2.
Working in $V_{0}$ we fix $j:V_{0}\rightarrow M_{0}$ such that $j$ witnesses that $\kappa$ is $\lambda_{n}$-supercompact, and $j(\mathbb{L})\simeq{\mathbb{L}}*{\mathbb{P}}*{\mathbb{R}}$ where the first point in the support of the tail iteration $\mathbb{R}$ is greater than $\lambda_{n}$. This ensures that:
@@
[…]
  2. (2)
    $V[G]\models{}^{\lambda_{n}}M_{0}[L][G]\subseteq M_{0}[L][G]$
  3. (3)
[…]
It is straightforward to lift $j:V_{0}\rightarrow M_{0}$ to obtain $j:V=V_{0}[L]\rightarrow M_{0}[L*G*H]$ for any $H$ which is $\mathbb{R}$-generic over $V[G]$. To further lift $j$ onto $V[G]$ we would need a $j({\mathbb{P}})$-generic object $S$ such that $j[G]\subseteq S$. It is worth noting that, since conditions in ${\mathbb{P}}$ have supports of size less than $\kappa$, the support of a condition in $j[G]$ will be contained in $j[I]$.
Let $\prec$ be the lexicographic ordering on $I$, so that:
  • $\prec$ is a well-ordering of $I$ with order type $\lambda_{n}$.
  • $\prec$ extends $<$.
Let $(a_{i})_{i<\lambda_{n}}$ enumerate $I$ in $\prec$-increasing order.
Working in $V[G]$, we will construct a decreasing sequence $(r_{i},s_{i})_{i<\lambda_{n}}$ of conditions in ${\mathbb{R}}*j(\mathbb{P})$. We will arrange that:
  • The support of $s_{i}$ is contained in $j(\{a_{\eta}:\eta<i\})$, so that in particular $j(a_{\zeta})$ is not in the support of $s_{i}$ for $\zeta\geq i$.
  • $s_{i}\leq j[G_{a_{i}}]$, so that $(r_{i},s_{i})$ forces that $j$ can be lifted onto $V[G_{a_{i}}]$.
We note that $s_{i}$ may not actually be a condition in $j({\mathbb{P}}_{a_{i}})$, the point is that the projection of $s_{i}$ to $j({\mathbb{P}}_{a_{i}})$ forces that the $j({\mathbb{P}}_{a_{i}})$-generic object contains $j[G_{a_{i}}]$.
For $i=0$, where we note that $a_{0}=(0,0,\ldots,0)$, let $(r_{0},s_{0})$ be the trivial condition. For $i$ limit let $(r_{i},s_{i})$ be a lower bound for $(r_{j},s_{j})_{j<i}$, taking care that the support of $s_{i}$ is the union of the supports of $s_{j}$ for $j<i$.
To construct $(r_{i+1},s_{i+1})$, we force below $(r_{i},s_{i})$ and construct $j:V[G_{a_{i}}]\rightarrow M_{0}[L][G][H][S^{\prime}]$ where $H*S^{\prime}$ is ${\mathbb{R}}*j({\mathbb{P}}_{a_{i}})$-generic. Recall that we defined ${\mathbb{Q}}_{a_{i}}$ from a $\kappa$-complete uniform filter $F_{a_{i}}\in V[G_{a_{i}}]$, and that the $a_{i}$-component of $G$ gives a subset $B_{a_{i}}$ of $\kappa$ which is ${\mathbb{Q}}_{a_{i}}$-generic.
Let $s^{\prime}_{i}$ be the projection (i.e. restriction) of $s_{i}$ to $j({\mathbb{P}}_{a_{i}})$. There are now two cases:
  1. (1)
    $(r_{i},s^{\prime}_{i})\Vdash^{M_{0}[L][G]}_{{\mathbb{R}}*j({\mathbb{P}}_{a_{i}})}\kappa\in j(E)$ for every $E\in F_{a_{i}}$. In this case we define $r_{i+1}$ and $s_{i+1}$ as follows:
@@
    There is $E\in F_{a_{i}}$ such that $(r_{i},s^{\prime}_{i})\not\Vdash^{M_{0}[L][G]}_{{\mathbb{R}}*j({\mathbb{P}}_{a_{i}})}\kappa\in j(E)$. In this case we choose such an $E_{i}=E$ and find $(r_{i}^{*},s_{i}^{*})\in{\mathbb{R}}*j({\mathbb{P}}_{a_{i}})$ such that $(r_{i}^{*},s_{i}^{*})\leq(r_{i},s^{\prime}_{i})$ and $(r_{i}^{*},s_{i}^{*})\Vdash\kappa\notin j(E_{i})$. We then define:
@@
      $\operatorname{dom}(s_{i+1})=\operatorname{dom}(s^{*}_{i})\cup(\operatorname{dom}(s_{i})\setminus\operatorname{dom}(s^{*}_{i}))\cup\{j(a_{i})\}$.
@@
    $j(F_{a_{i}})$ is $j(\kappa)$-complete and $j[F_{a_{i}}]\in M_{0}[L][G][H][S^{\prime}]$ so that $\bigcap j[F_{a_{i}}]\in j(F_{a_{i}})$ and we constructed legitimate conditions in $j({\mathbb{Q}}_{a_{i}})$.
@@
    In either case of the construction, $\operatorname{dom}(s_{i+1})\subseteq\operatorname{dom}(s_{i})\cup j({I_{<a_{i}}})\cup\{j(a_{i})\}$. Since $I_{<a_{i}}\subseteq\{a_{\eta}:\eta<i\}$, it follows that $\operatorname{dom}(s_{i+1})\subseteq j(\{a_{\zeta}:\zeta\leq i\})$.
@@
    The filter added by $G$ at $a_{i}$ is the set of pairs $(s,E)$ where $E\in F_{a}$, $s$ is an initial segment of $B_{a_{i}}$ and $B_{a_{i}}\setminus s\subseteq E$. Since $j((s,E))=(s,j(E))$, it is easy to see that it is forced that $(B_{a_{i}},\bigcap j[F_{a_{i}}])\leq(s,j(E))$. In the first case $(r_{i},s_{i})$ forces that $(B_{a_{i}}\cup\{\kappa\},\bigcap j[F_{a_{i}}])\leq(s,j(E))$.
When the construction of $(r_{i},s_{i})_{i<\lambda_{n}}$ is done, we choose $(r^{\prime},s^{\prime})$ to be a lower bound and note that $(r^{\prime},s^{\prime})$ forces that $j$ can be lifted onto $V[G]$. Recall that every subset of $\kappa$ in $V[G]$ lies in $V[G_{a}]$ for some $a\in I$, and observe that if $X=\{\dot{A}:\mbox{For some $a\in I$, $\dot{A}$ is a canonical ${\mathbb{P}}_{a}$-name for a subset of $\kappa$}\}$ then $|X|=\lambda_{n}$. Using the closure of ${\mathbb{R}}*j(\mathbb{P})$ again, we find $(r,s)\leq(r^{\prime},s^{\prime})$ such that $(r,s)$ decides $\kappa\in j(\dot{A})$ for every $\dot{A}\in X$. It is now easy to see that if $U=\{{\dot{A}}[G]:\mbox{$\dot{A}\in X$ and $(r,s)\Vdash\kappa\in j(\dot{A})$}\}$ then $V[G]\models\mbox{``$U$ is a measure on $\kappa$''}$ ”.
Let $\dot{U}$ be a canonical name for the measure $U$ produced by this construction, and for $\beta<\lambda_{n}$ let
$\dot{U}_{\beta}=\{(q,\dot{\tau})\in\dot{U}:$ $q\in{\mathbb{P}}_{\prod_{i<n}\lambda_{i}\times\beta}\text{ and}$
$\dot{\tau}\text{ is a canonical }{\mathbb{P}}_{\prod_{i<n}\lambda_{i}\times\beta}\text{-name for a subset of }\kappa\}$
Then:
  • $\dot{U}\subseteq H_{\lambda_{n}}$.
  • $\dot{U}=\bigcup_{\beta<\lambda_{n}}{\dot{U}}_{\beta}$, and $\langle\dot{U}_{\beta}\mid\beta<\lambda\rangle$ is continuous at points of cofinality $\kappa^{+}$.
  • For almost every $\beta\in\lambda_{n}\cap\operatorname{cof}(\kappa^{+})$, $\dot{U}_{\beta}$ is a name for a measure on $\kappa$, and is forced to equal $U\cap V[G_{\prod_{i<n}\lambda_{i}\times\beta}]$.
  • If $S=\{\beta\in\lambda_{n}\cap\operatorname{cof}(\kappa^{+}):A_{\beta}=\dot{U}_{\beta}\}$ then $S$ is stationary in $\lambda_{n}$, and every $\beta\in S$ is an “active stage” in the definition of $\mathbb{P}$.
Let $I^{*}=\lambda_{0}\times\ldots\lambda_{n-1}\times S$, and note that $(I^{*},<)$ is isomorphic to $(I,<)$.
@@
Let $a=a_{i}=(\alpha_{0},\ldots\alpha_{n-1},\beta)$, and consider the construction of $(r_{i+1},s_{i+1})$. The proof is by induction on $a$. The key point is that we must be in case 1 of the construction, so assume for a contradiction that we are in case 2. Since $(r,s)\leq(r_{i+1},s_{i+1})$, this means that there is $E\in F_{a_{i}}$ such that $E\notin U$. Since $U$ is $\kappa$-complete and since $F_{a_{i}}$ is the $\kappa$-complete filter generated by $F^{1}_{a}\cup\mathcal{X}_{a}$, either there is $Y\in A_{\beta}[G_{\lambda_{0}\times\ldots\lambda_{n-1}\times\beta}]\setminus U$ or $B_{a^{\prime}}\notin U$ for some $a^{\prime}=(\alpha_{1}^{\prime},...,\alpha_{n-1}^{\prime},\beta)<a$. Note that $a^{\prime}\in I^{*}$ as well, hence, by the induction hypothesis, $B_{a^{\prime}}\in U$. In the other case, since $\beta$ is an active stage, $A_{\beta}=\dot{U}_{\beta}$, and thus $Y\in U\cap V[G_{\prod_{i<n}\lambda_{i}\times\beta}]$. In either case, we reach a contradiction.
@@
As a corollary, we conclude that whenever $\beta$ is an active stage of the iteration, for every $a=(\alpha_{1},...,\alpha_{n-1},\beta)$, $F^{1}_{a}\cup\mathcal{X}_{a}\subseteq U$ and therefore generates a $\kappa$-complete ultrafilter. By the definition of $F_{a}$, this means that $F_{1}\cup\mathcal{X}_{a}\subseteq F_{a}$.
@@
Let $a,a^{\prime}\in I^{*}$ where $a<a^{\prime}$. Then $B_{a}\in F_{a^{\prime}}$.

### v2

@@ 4.1. At a measurable cardinal
This problem was first solved by Džamonja and Shelah [21], where the rough idea is that we force at $\alpha$ with the disjoint sum of all measures which exist in $V[G_{\alpha}]$. This leads to some technical complications in the iteration, which we can avoid using the diamond machinery we used to define the iteration which traces back to Magidor. The key feature of the construction of $U$ is to ensure that whenever $U_{\alpha}\subseteq U$ (which happens frequently) we have $B_{(\alpha,d)}\in U$ for every $d\in\mathbb{D}$. Running the iteration for a large enough number of steps, we may produce measures which are generated by arbitrarily long $\subseteq^{*}$-decreasing sequences. The diamond idea appears in a paper by Cummings, Džamonja, Magidor, Morgan and Shelah [18] albeit in a more complex setting.
Moving on towards the proof of 4.6, we start with a model $V_{0}$ in which $\kappa$ is supercompact and $\mathbb{D}$ is a suitable directed set with a linearisation $\langle d_{\alpha}:\alpha<\mu\rangle$. We start by forcing with the Laver preparation $\mathbb{L}$ from [35] which makes $\kappa$ indestructibly supercompact. Next, we let $\mathbb{Q}$ be the standard forcing to add a $\diamondsuit(\mu^{+}\cap cf(\mu))$-sequence by initial segments. We note that this forcing is $\mu^{+}$-directed closed and if $2^{\mu}>\mu^{+}$, it collapses $2^{\mu}$ to $\mu^{+}$. We obtain a model
$V\models$ $``\kappa\mbox{ is supercompact}\mbox{ indestructible under }$
$\kappa\mbox{-directed closed forcing}+\diamondsuit(\mu^{+}\cap\operatorname{cof}(\mu))$".
Working in $V$ we let $\mathbb{P}={\mathbb{P}}_{{\mathbb{D}}^{+}}$.
@@
If $G$ is $\mathbb{P}$-generic then there exist in $V[G]$ a normal measure $U$ on $\kappa$, such that a canonical name $\dot{U}$ for $U$ will have the set $S^{*}_{\dot{U}}\times\mathbb{D}$ satisfy ($\dagger$3.5).
Proof of Theorem 4.7.
Working in $V_{0}$ we fix $j:V_{0}\rightarrow M_{0}$ such that $j$ witnesses that $\kappa$ is $\mu^{+}$-supercompact and $j(\mathbb{L})\simeq{\mathbb{L}}*{\mathbb{P}}*{\mathbb{R}}$ where the first point in the support of the tail iteration $\mathbb{R}$ is greater than $\mu^{+}$. This ensures that:
@@
    $\Vdash^{M_{0}}_{{\mathbb{L}}*{\mathbb{P}}}\mbox{``$\mathbb{R}*j(\mathbb{P})$ is $\mu^{++}$-closed''}$.
@@
    $V[G]\models\mbox{``$\mathbb{R}*j(\mathbb{P})$ is $\mu^{++}$-closed''}$.
It is straightforward to lift $j:V_{0}\rightarrow M_{0}$ to obtain $j:V=V_{0}[L]\rightarrow M_{0}[L*G*H]$ for any $H$ which is $\mathbb{R}$-generic over $V[G]$. To further lift $j$ onto $V[G]$ we would need a $j({\mathbb{P}})$-generic object $S$ such that $j[G]\subseteq S$. It is worth noting that, since conditions in $\mathbb{P}$} have supports of size less than $\kappa$, the support of a condition in $j[G]$ will be contained in $j[\mathbb{D}^{+}]$.
Let $\prec$ be the “lexicographic” ordering of $\mathbb{D}^{+}$ given by $(\epsilon,d_{\alpha})\prec(\zeta,d_{\beta})$ if and only if eithe r $\epsilon<\zeta$ or $\epsilon=\zeta$ and $\alpha<\beta$. It is easy to see that:
  • $\prec$ is a well-ordering of $\mathbb{D}^{+}$ with order type $\mu^{+}$.
  • $\prec$ extends $_{{\mathbb{D}}^{+}}$.
Denote by $\{e_{\zeta}\mid\zeta<\mu^{+}\}$ the enumeration of $\mathbb{D}^{+}$ in $\prec$-increasing order.
Working in $V[G]$, we will construct a decreasing sequence $(r_{i},s_{i})_{i<\mu^{+}}$ of conditions in ${\mathbb{R}}*j(\mathbb{P})$. We will arrange that:
  • The support of $s_{i}$ is contained in $j(\{e_{\eta}:\eta<i\})$, so that in particular $j(e_{\zeta})$ is not in the support of $s_{i}$ for $\zeta\geq i$.
  • $s_{i}\leq j[G_{<e_{i}}]$, so that $(r_{i},s_{i})$ forces that $j$ can be lifted onto $V[G_{<e_{i}}]$.
We note that $s_{i}$ may not actually be a condition in $j({\mathbb{P}}_{<e_{i}})$, the point is that the projection of $s_{i}$ to $j({\mathbb{P}}_{<e_{i}})$ forces that the $j({\mathbb{P}}_{<e_{i}})$-generic object contains $j[G_{<e_{i}}]$.
For $i=0$, there is no element of $\mathbb{D}^{+}$ below $e_{0}$, so let $(r_{0},s_{0})$ be the trivial condition. For $i$ non zero limit ordinal let $(r_{i},s_{i})$ be a lower bound for $(r_{j},s_{j})_{j<i}$, taking care that the support of $s_{i}$ is the union of the supports of $s_{j}$ for $j<i$.
To construct $(r_{i+1},s_{i+1})$, we force below $(r_{i},s_{i})$ and construct $j:V[G_{<e_{i}}]\rightarrow M_{0}[L][G][H][S^{\prime}]$ where $H*S^{\prime}$ is ${\mathbb{R}}*j({\mathbb{P}}_{<e_{i}})$-generic. Recall that we defined ${\mathbb{Q}}_{e_{i}}$ from a $\kappa$-complete uniform filter $F_{e_{i}}\in V[G_{<e_{i}}]$, and that the $e_{i}$-component of $G$ gives a subset $B_{e_{i}}$ of $\kappa$ which is ${\mathbb{Q}}_{e_{i}}$-generic.
Let $s^{\prime}_{i}$ be the projection (i.e. restriction) of $s_{i}$ to $j({\mathbb{P}}_{<e_{i}})$. There are now two cases:
@@
    $(r_{i},s^{\prime}_{i})\Vdash^{M_{0}[L][G]}_{{\mathbb{R}}*j({\mathbb{P}}_{<e_{i}})}\kappa\in j(E)$ for every $E\in F_{e_{i}}$. In this case we define $r_{i+1}$ and $s_{i+1}$ as follows:
@@
    There is $E\in F_{e_{i}}$ such that $(r_{i},s^{\prime}_{i})\not\Vdash^{M_{0}[L][G]}_{{\mathbb{R}}*j({\mathbb{P}}_{<e_{i}})}\kappa\in j(E)$. In this case we choose such an $E_{i}=E$ and find $(r_{i}^{*},s_{i}^{*})\in{\mathbb{R}}*j({\mathbb{P}}_{<e_{i}})$ such that $(r_{i}^{*},s_{i}^{*})\leq(r_{i},s^{\prime}_{i})$ and $(r_{i}^{*},s_{i}^{*})\Vdash\kappa\notin j(E_{i})$. We then define:
@@
      $\operatorname{dom}(s_{i+1})=\operatorname{dom}(s^{*}_{i})\cup(\operatorname{dom}(s_{i})\setminus\operatorname{dom}(s^{*}_{i}))\cup\{j(e_{i})\}$.
@@
    $j(F_{e_{i}})$ is $j(\kappa)$-complete and $j[F_{e_{i}}]\in M_{0}[L][G][H][S^{\prime}]$ so that $\bigcap j[F_{e_{i}}]\in j(F_{e_{i}})$ and we constructed legitimate conditions in $j({\mathbb{Q}}_{e_{i}})$.
@@
    In either case of the construction, $\operatorname{dom}(s_{i+1})\subseteq\operatorname{dom}(s_{i})\cup j({\mathbb{D}_{<e_{i}}})\cup\{j(e_{i})\}$. Since $\mathbb{D}_{<e_{i}}\subseteq\{e_{\eta}:\eta<i\}$, it follows that $\operatorname{dom}(s_{i+1})\subseteq j(\{e_{\zeta}:\zeta\leq i\})$.
@@
    The filter added by $G$ at $e_{i}$ is the set of pairs $(s,E)$ where $E\in F_{e_{i}}$, $s$ is an initial segment of $B_{e_{i}}$ and $B_{e_{i}}\setminus s\subseteq E$. Since $j((s,E))=(s,j(E))$, it is easy to see that it is forced that $(B_{e_{i}},\bigcap j[F_{e_{i}}])\leq(s,j(E))$. In the first case $(r_{i},s_{i})$ forces that $(B_{e_{i}}\cup\{\kappa\},\bigcap j[F_{e_{i}}])\leq(s,j(E))$.
When the construction of $(r_{i},s_{i})_{i<\mu^{+}}$ is done, we choose $(r^{\prime},s^{\prime})$ to be a lower bound and note that $(r^{\prime},s^{\prime})$ forces that $j$ can be lifted onto $V[G]$. Recall that every subset of $\kappa$ in $V[G]$ lies in $V[G_{<a}]$ for some $a\in\mathbb{D}^{+}$, and observe that if
[…]
then $|X|=\mu^{+}$. Using the closure of ${\mathbb{R}}*j(\mathbb{P})$ again, we find $(r,s)\leq(r^{\prime},s^{\prime})$ such that $(r,s)$ decides $\kappa\in j(\dot{A})$ for every $\dot{A}\in X$. It is now easy to see that if $U=\{{\dot{A}}[G]:\mbox{$\dot{A}\in X$ and $(r,s)\Vdash\kappa\in j(\dot{A})$}\}$ then $V[G]\models\mbox{``$U$ is a measure on $\kappa$''}$.
Let $\dot{U}$ be a canonical name for the measure $U$ produced by this construction and let us verify ($\dagger$3.5) of Corollary 3.5. This is done in Claim 4.9:
@@
Let $a=e_{i}=(\beta,d)$ say, and consider the construction of $(r_{i+1},s_{i+1})$. The key point is that we must be in case 1 of the construction, so assume for a contradiction that we are in case 2. Since $(r,s)\leq(r_{i+1},s_{i+1})$, this means that there is $E\in F_{e_{i}}$ such that $E\notin U$. Since $\beta$ is an active stage, $A_{\beta}=\dot{U}_{\beta}$, and thus $E\in U\cap V[G_{\beta\times\mathbb{D}}]$, which is a contradiction.
@@
As a corollary, we conclude that whenever $\beta$ is an active stage of the iteration the sets $B_{a}$ for $a\in\{\beta\}\times\mathbb{D}$ belong to $U$, and therefore generate a $\kappa$-complete filter.
@@
Let $a,a^{\prime}\in S^{*}_{\dot{U}}\times\mathbb{D}$ where $a<a^{\prime}$. Then $\Vdash_{\mathbb{P}_{<a^{\prime}}}\dot{B}_{a}\in\dot{F}_{a^{\prime}}$.